# Algebraically Skew Embedding of Curves

### v1

@@
Abstract: Given a smooth complex variety $X$, an algebraically skew embedding of $X$ is an embedding of $X$ into a complex projective space $\mathbb{P}^{N}$ such that for any two points $x,y\in X$, their embedded tangent spaces in $\mathbb{P}^{N}$ do not intersect. In this work, we establish an upper bound and a lower bound of the minimal dimension $N=\mathsf{msdim}\,X$ such that there exists an algebraically skew embedding into $\mathbb{P}^{N}$ in terms of the dimension of the given smooth variety $X$. Then we further classify the algebraic curves in terms of their minimal skew embedding dimensions.
1 Introduction
History and motivations
The questions of skew embedding started in a lecture given by H. Steinhaus in 1966. In the lecture, he asked: Are there smooth loops in the space without a pair of parallel tangents? He called such a curve a skew loop. The problem was soon solved by B. Segre in [Seg68] in 1968. Segre used convex geometry to construct several skew loops in $\mathbb{R}^{3}$. Despite the existence of these skew loops, Segre noted that such loops lie on neither ellipsoids nor elliptic paraboloids. Later, Q-Y. Wu showed that there exist skew loops of every knot class [Wu02], and B. Solomon further showed that skew loops of every knot class also exist in the flat $3$-torus [Sol07]. On the other hand, M. Ghomi and Solomon showed that Ellipsoids are the only closed surfaces that contain no skew loops [GS02].
@@ 1 Introduction
Meanwhile, in [GT08], Ghomi and Tabachnikov asked: Given any manifold $M$, does there exist an embedding of $M$ to an euclidean space $\mathbb{R}^{N}$ such that for each pair of points in the image of $M$, their tangent spaces to $M$ neither intersect nor contain parallel lines? If so, what is the minimum dimension of $\mathbb{R}^{N}$? Ghomi and Tabachnikov called such an embedding a totally skew embedding. In [GT08], Tabachnikov proved, among other things, that such an embedding always exists, and for any $n$-dimensional manifold $M$, the minimum dimension of $\mathbb{R}^{N}$ is always between $12n+$ 1 and $14n+$ 1.
On the one hand, the condition of the totally skewness defined above is equivalent to saying that for each pair of points of the image of $M$, their tangent spaces do not intersect in the projectification of $\mathbb{R}^{N}$. On the other hand, we know that there is a rich theory in algebraic geometry on the intersection of subvarieties in a projective space, especially for spaces over an algebraically closed characteristic zero field. Thus, it is natural to study the skew embeddings of complex projective varieties in complex projective spaces.
@@
For any subset $S\subseteq\mathbb{P}^{N}$, we denote its Zariski closure by $\overline{S}$. For any variety $X$, we denote its tangent sheaf (or tangent bundle if $X$ is smooth) by $\mathcal{T}_{X}$ and the tangent space at a smooth point $x\in X$ by $\mathcal{T}_{x,X}$. For a pair of smooth varieties $X\subseteq Y$, we write $\mathcal{N}_{X,Y}$ for the normal bundle of $X$ over $Y$ and $\mathcal{N}_{x,X,Y}$ for the fibre of $\mathcal{N}_{X,Y}$ at $x$. For any projective subvariety $X\subseteq\mathbb{P}$ and any smooth point $x\in X$, we write $T_{x}=T_{x,X}\subseteq\mathbb{P}$ for the embedded tangent space of $X$ at $x$. For any two planes (linear subspaces) $K,L\subseteq\mathbb{P}$, we use $\left\langle K,L\right\rangle$ to denote the plane spanned by $K$ and $L$ in $\mathbb{P}$.
@@
For a general projective variety $X\subseteq\mathbb{P}$, one possible definition for the algebraically skewness is to require that the projective second and third fundamental forms being nowhere vanishing. The nowhere vanishing condition of the second projective fundamental form is equivalent to say that the Gauss map from $X$ to the appropriate Grassmannian is a diffeomorphism to its image, and the nowhere vanishing condition of the third projective fundamental form would exclude the infinitesimal nonskewness. For more on the projective fundamental forms, see [IL03].
@@
For any projective variety $X$, we denote the minimum algebraically skew embedding dimension of $X$ by
@@
In Section 2, we consider the dimension of the tangent variety of the secant variety of a projective variety and use it to deduce that $\mathsf{msdim}\,X\leq 4n+1$ for any $n$-dimensional variety $X$. Then we introduce the degeneracy loci $\mathbb{D}_{r}$ of pairs of linear subspaces in $\mathbb{P}$ with intersecting dimension $\geq r$ in ${\mathbb{G}\mathrm{r}}(n,N)\times{\mathbb{G}\mathrm{r}}(n,N)$. By considering the dimension of the intersection between $D_{1}=\mathbb{D}_{0}$ and $\Gamma:=\gamma(X)\times\gamma(X)$, the product of the Gauss image of $X$, we deduce that $3n\leq\mathsf{msdim}\,X$. Note that the upper bound $14n+$ 1 agrees with the real smooth case, but the lower bound $3n$ is better than the real smooth case, which is $12n+$ 1.
@@
Although the intersection $D_{1}\cap\Gamma$ we study in Section 2 gives us a lower bound of $\mathsf{msdim}\,X$, it is not enough to solve $\mathsf{msdim}\,X$ for a specific variety $X$ because $D_{1}$ is singular on the diagonal of $G\times G$. We resolve the singularity of $D_{1}$ for the case of curves in Section 3 by blowing up the diagonal of $G\times G$. Then we compute the intersection between the proper transforms $\widetilde{D}_{1}$ and $\widetilde{\Gamma}$ of $D_{1}$ and $\Gamma$ for the case of $X\subseteq\mathbb{P}^{3}$. By comparing the product $[\widetilde{D}_{1}][\widetilde{\Gamma}]$ and the class of the diagonal component $\widetilde{\Delta}_{\Gamma}$ of the intersection in the Chow ring of the blowup $B:=\mathsf{Bl}_{\Delta_{G}}(G\times G)$, we conclude that the twisted cubic is the only skew curve in $\mathbb{P}^{3}$.
@@
In Section 4, we study the case of curves in $\mathbb{P}^{4}$. By dimension counting we immediately see that $\widetilde{D}_{1}$ and $\widetilde{\Gamma}$ have an excess intersection component at the diagonal component $\widetilde{\Delta}_{\Gamma}$. In addition, we discover that $\widetilde{D}_{1}$ and $\widetilde{\Gamma}$ in fact tangent along $\widetilde{\Delta}_{\Gamma}$. We resolve this tangency by blowing up $B:=\mathsf{Bl}_{\Delta_{G}}(G\times G)$ along $\widetilde{\Delta}_{\Gamma}$. Then we consider the intersection class between the proper transforms $D_{1}^{\dagger}$ and $\Gamma^{\dagger}$ of $\widetilde{D}_{1}$ and $\widetilde{\Gamma}$ as well as the excess intersection class $\varepsilon_{\Delta^{\dagger}_{\Gamma}}$ of the component on the diagonal. The main obstruction of computing the excess intersection class in question is the Chow ring of $\widetilde{D}_{1}$ and the class of $\widetilde{\Delta}_{\Gamma}$ in it. We establish the structure of $\mathsf{A}(\widetilde{D}_{1})$ by introducing an intermediate incidence variety $\hat{D}_{1}\subseteq\mathbb{P}\times G\times G$ between $\widetilde{D}_{1}$ and $D_{1}$. We discover that $\hat{D}_{1}$ can be treated as a biprojective bundle $\mathbb{P}(\mathcal{Q})\times_{\mathbb{P}^{4}}\mathbb{P}(\mathcal{Q})$. In addition, we interpret $\widetilde{D}_{1}$ as the blowup of $\hat{D}_{1}$ along the fibrewise diagonal $\hat{\Delta}_{D}$. We establish the structure $\mathsf{A}(\widetilde{D}_{1})$ by realizing that $\hat{\Delta}_{D}$ is isomorphic a flag variety and the normal bundle of $\hat{\Delta}_{D}$ in $\hat{D}_{1}$ is isomorphic to a subbundle of $\mathcal{T}_{\hat{\Delta}_{D}}$. With these tools, we manage to compute the excess intersection term. Then we assert that a curve $X\subseteq\mathbb{P}^{4}$ is algebraically skew if and only if the intersection class $[D_{1}^{\dagger}][\Gamma^{\dagger}]$ agrees with the excess intersection class $\varepsilon_{\Delta^{\dagger}_{\Gamma}}$ from the diagonal and conclude that the only algebraically skew curves in $\mathbb{P}^{4}$ are the rational normal curve and the elliptic normal curves.
@@ Acknowledgment
The author is a graduate student at Pennsylvania State University. The author is grateful to his advisors Sergei Tabachnikov and John Lesieutre for the weekly meetings in writing this paper. The author thanks Jack Huizenga for several crucial conversations and advice. The author also benefits from several insightful email exchanges with Joseph Landsberg and Serge Lvovski. The author also thanks Brendan Hassett for pointing out the potential double point formula approach to the problem. The author also thanks Dawei Chen for a discussion on whether the intersection component on the blown-up diagonal in the present paper is a ribbon.
2 The basic bounds of $\mathsf{msdim}\,X$
In this section, we prove the existence of skew embeddings of a projective variety $X$ by showing that the Veronese embeddings of degree $\geq 3$ are skew and establish the general upper bound of the minimum complex algebraic skew embedding dimension $\mathsf{msdim}\,X$ by investigating the tangent of the secant variety. Then we introduce the degeneracy filtration of $\mathsf{Gr}(n,N)\times\mathsf{Gr}(n,N)$, and use them to compute a lower bound of $\mathsf{msdim}\,X$.
2.1 Skewness of higher degree Veronese embeddings
Following the same approach for the real smooth case, one can easily show that the twisted cubic is a skew embedding. This result can be generalized to all Veronese curves of degree $\geq 3$.
@@ 2.1 Skewness of higher degree Veronese embeddings
Proof.
It is known that the map $\iota_{d}$ is an embedding. To verify the skewness, we first consider the affine chart $s\neq 0$. Set $u:=t/s$, we obtain $\iota_{d}(u)=(u,u^{2},\cdots,u^{d-1})$. Project this to the first three coordinates yields $u\longmapsto(u,u^{2},u^{3})$. This is the standard twisted cubic, which is skew. Clearly, if the projection of a curve is skew, then the curve itself is skew.
Similarly, $\iota_{d}$ restricted to $t\neq 0$ is also skew. So, it remains to show that the tangent lines of $\iota_{d}$ at $\iota_{d}(0:1)$ and $\iota_{d}(1:0)$ do not meet each other. This can be verified by straightforward computation. ∎
By the observation above, one can prove that all Veronese maps of degree $\geq 3$ are skew.
@@
$\alpha\widetilde{\iota}_{d}(\widetilde{x})+\sum_{k=0}^{n}a_{k}\frac{\partial\widetilde{\iota}_{d}}{\partial x_{k}}(\widetilde{x})=\beta\widetilde{\iota}_{d}(\widetilde{y})+\sum_{k=0}^{n}b_{k}\frac{\partial\widetilde{\iota}_{d}}{\partial x_{k}}(\widetilde{y}).$ (1)
@@
2.2 The basic upper bound of $\mathsf{msdim}\,X$
To find the basic upper bound of $\mathsf{msdim}\,X$, suppose that $X\hookrightarrow\mathbb{P}^{N}$ is a skew embedding with $N=\mathsf{msdim}\,X$. Consider the secant variety $\mathsf{Sec}\,X\subseteq\mathbb{P}^{N}$ of $X$. From [Zak05, Proposition II.1.10], we have

### v4

# Algebraically Skew Embeddings of Curves

@@
Abstract.
Given a smooth complex variety $X$, an algebraically skew embedding of $X$ is an embedding of $X$ into a complex projective space $\mathbb{P}^{N}$ such that for any two points $x,y\in X$, their embedded tangent spaces in $\mathbb{P}^{N}$ do not intersect. In this work, we establish an upper bound and a lower bound of the minimal dimension $N=\mathsf{msdim}\,X$ such that there exists an algebraically skew embedding into $\mathbb{P}^{N}$ in terms of the dimension of the given smooth variety $X$. Then we further classify the algebraic curves in terms of their minimal skew embedding dimensions, and apply the same technique to other one-parameter family of lines.
Key words and phrases:
skew embedding, algebraic curves, secant variety, Terracini loci
1. Introduction
History, motivations, and relevant works
This work is motivated by the skew embedding problems from differential topology. The questions of skew embedding started in a lecture given by H. Steinhaus in 1966. In the lecture, he asked: Are there smooth loops in the space without a pair of parallel tangents? He called such a curve a skew loop. The problem was soon solved by B. Segre in [Seg68] in 1968. Segre used convex geometry to construct several skew loops in $\mathbb{R}^{3}$. Despite the existence of these skew loops, Segre noted that such loops lie on neither ellipsoids nor elliptic paraboloids. Later, Q-Y. Wu showed that there exist skew loops of every knot class [Wu02], and B. Solomon further showed that skew loops of every knot class also exist in the flat $3$-torus [Sol07]. On the other hand, M. Ghomi and Solomon showed that Ellipsoids are the only closed surfaces that contain no skew loops [GS02].
@@ 1. Introduction
Meanwhile, in [GT08], Ghomi and Tabachnikov asked: Given any manifold $M$, does there exist an embedding of $M$ to an euclidean space $\mathbb{R}^{N}$ such that for each pair of points in the image of $M$, their tangent spaces to $M$ neither intersect nor contain parallel lines? Ghomi and Tabachnikov called such an embedding a totally skew embedding. They further asked, if such embeddings exist, what is the minimal dimension of $\mathbb{R}^{N}$? In [GT08], Tabachnikov proved, among other things, that such an embedding always exists, and for any $n$-dimensional manifold $M$, the minimal dimension of $\mathbb{R}^{N}$ is always between $12n+$ 1 and $14n+$ 1.
On the one hand, the condition of the totally skewness defined above is equivalent to saying that for each pair of points of the image of $M$, their tangent spaces do not intersect in the projectification of $\mathbb{R}^{N}$. On the other hand, we know that there is a rich theory in algebraic geometry on the intersection of subvarieties in a projective space, especially for spaces over an algebraically closed characteristic zero field. Thus, it is natural to study the skew embeddings of complex projective varieties in complex projective spaces. In this work, we take on the Gauss map approach for the complex projective subvarieties, and investigate other applications of the method we developed.
It turns out that the complex projective version of skew embeddings fits into the theory around Terracini loci nicely. The notion of Terracini loci was first introduced in [BC21] by E. Ballico and L. Chiantini. Roughly speaking, the Terracini loci of a smooth projective subvariety $X\subseteq\mathbb{P}^{N}$ capture the finite subsets of $X$ such that the tangent spaces at the points in a given set are not in general position. The Terracini loci of Veronese embeddings have been studied in [BC21] and [BV23], and the ones of Segre embeddings have been studied in [BBS23]. In [BC23] E. Ballico and L. Chiantini have looked into the Terracini loci of the curves in $\mathbb{P}^{3}$ and of the canonical curves.
C. Ciliberto kindly informed us that in his recent paper [Cil24], he studied the Terracini loci of curves in $\mathbb{P}^{4}$, which coincides with Lemma 4.11 in the present paper (while Lemma 4.11 counts the number of “non-skew” ordered pairs of tangent lines instead of the length of the Terracini locus). Our approach to proving Lemma 4.11 is based on an explicit blowup construction, which differs from the one in Ciliberto’s work. In [Cil24], Ciliberto observed that, by pulling back the theta divisor to the incidence variety associated to the Terracini locus, one can apply the Grothendieck-Riemann-Roch theorem and the Porteous formula to compute the length of the Terracini loci. We believe our method suggest new directions different from those in the existing Terracini loci literature. For instance, our method can be applied to calculate the locus of intersection of families of linear projective subspaces, as described in Section 3.5 and Section 4.2. More significantly, the difficulty encountered in the surface case discussed in [Cil24] does not appear to arise with our approach. Thus, the present paper offers a different approach for studying the Terracini loci of higher-dimensional varieties.
Independently, the complex version of affine spacial skew curves has occurred in [Zai99].
@@
For any subset $U\subseteq\mathbb{P}^{N}$, we denote its Zariski closure by $\overline{U}$. For any variety $X$, we denote its tangent sheaf (or tangent bundle if $X$ is smooth) by $\mathcal{T}_{X}$ and the tangent space at a smooth point $x\in X$ by $\mathcal{T}_{x,X}$. For a pair of smooth varieties $X\subseteq Y$, we write $\mathcal{N}_{X,Y}$ for the normal bundle of $X$ over $Y$ and $\mathcal{N}_{x,X,Y}$ for the fibre of $\mathcal{N}_{X,Y}$ at $x$. For any projective subvariety $X\subseteq\mathbb{P}$ and any smooth point $x\in X$, we write $T_{x}=T_{x,X}\subseteq\mathbb{P}$ for the embedded tangent space of $X$ at $x$. For any two planes (linear subspaces) $K,L\subseteq\mathbb{P}$, we use $\left\langle K,L\right\rangle$ to denote the plane spanned by $K$ and $L$ in $\mathbb{P}$.
@@
For a general projective variety $X\subseteq\mathbb{P}$, there are several possible definitions for algebraically skewness. One such definition is to define the algebraically skewness of $X\subseteq\mathbb{P}^{N}=\mathbb{P}(V)$ by vanishing of the length $2$ Terracini locus $\mathbb{T}(X,V,2)$, which consists of length $2$ zero dimensional subschemes $S\subseteq X$ satisfying $\mathsf{dim}\,V(-2S)>\mathsf{dim}\,V-2n-2$, c.f. [BC23].
Another possible definition for algebraically skewness is to require that the projective second and third fundamental forms being nowhere vanishing. The nowhere vanishing condition of the second projective fundamental form is equivalent to say that the Gauss map from $X$ to the appropriate Grassmannian is a diffeomorphism to its image, and the nowhere vanishing condition of the third projective fundamental form would exclude the infinitesimal nonskewness. For more on the projective fundamental forms, see [IL03].
@@
For any projective variety $X$, we denote the minimal algebraically skew embedding dimension of $X$ by
@@
In Section 2, we consider the dimension of the tangent variety of the secant variety of a projective variety and use it to deduce that $\mathsf{msdim}\,X\leq 4n+1$ for any $n$-dimensional variety $X$. Then we introduce the degeneracy loci $\mathbb{D}_{r}$ of pairs of linear subspaces in $\mathbb{P}$ with intersecting dimension $\geq r$ in ${\mathbb{G}\mathrm{r}}(n,N)\times{\mathbb{G}\mathrm{r}}(n,N)$. By considering the dimension of the intersection between $D_{1}=\mathbb{D}_{0}$ and $\Gamma:=C\times C\subseteq{\mathbb{G}\mathrm{r}}(n,N)\times{\mathbb{G}\mathrm{r}}(n,N)$, the product of the Gauss image $C:=\gamma(X)$ of $X$, we deduce that $3n\leq\mathsf{msdim}\,X$. Note that the upper bound $14n+$ 1 agrees with the real smooth case, but the lower bound $3n$ is better than the real smooth case, which is $12n+$ 1.
@@
Although the intersection $D_{1}\cap\Gamma$ we study in Section 2 gives us a lower bound of $\mathsf{msdim}\,X$, it is not enough to solve $\mathsf{msdim}\,X$ for a specific variety $X$ because $D_{1}$ is singular on the diagonal of $G\times G$. We resolve the singularity of $D_{1}$ for the case of curves in Section 3 by blowing up the diagonal of $G\times G$. Then we compute the intersection between the proper transforms $\widetilde{D}_{1}$ and $\widetilde{\Gamma}$ of $D_{1}$ and $\Gamma$ for the case of $X\subseteq\mathbb{P}^{3}$. By comparing the product $[\widetilde{D}_{1}][\widetilde{\Gamma}]$ and the class of the diagonal component $\widetilde{\Delta}_{\Gamma}$ of the intersection in the Chow ring of the blowup $B:=\mathsf{Bl}_{\Delta_{G}}(G\times G)$, we conclude that the twisted cubic is the only skew curve in $\mathbb{P}^{3}$. We also apply the same method to other one-parameter families $C\subseteq G$ and deduce that for a scroll to be skew, we must have $C\cong\mathbb{P}^{1}$.
@@
In Section 4, we study the case of one-parameter families of lines in $\mathbb{P}^{4}$. By dimension counting we immediately see that $\widetilde{D}_{1}$ and $\widetilde{\Gamma}$ have an excess intersection component at the diagonal component $\widetilde{\Delta}_{\Gamma}$. While the same method in section 3 works for a generic one-parameter family of lines, we discover that $\widetilde{D}_{1}$ and $\widetilde{\Gamma}$ tangent along $\widetilde{\Delta}_{\Gamma}$ when $\Gamma$ is the product of Gauss images. We resolve this tangency by blowing up $B:=\mathsf{Bl}_{\Delta_{G}}(G\times G)$ along $\widetilde{\Delta}_{\Gamma}$. Then we consider the intersection class between the proper transforms $D_{1}^{\dagger}$ and $\Gamma^{\dagger}$ of $\widetilde{D}_{1}$ and $\widetilde{\Gamma}$ as well as the excess intersection class $\varepsilon_{\Delta^{\dagger}_{\Gamma}}$ of the component on the diagonal. The main obstruction of computing the excess intersection class in question is the Chow ring of $\widetilde{D}_{1}$ and the class of $\widetilde{\Delta}_{\Gamma}$ in it. We establish the structure of $\mathsf{A}(\widetilde{D}_{1})$ by introducing an intermediate incidence variety $\hat{D}_{1}\subseteq\mathbb{P}\times G\times G$ between $\widetilde{D}_{1}$ and $D_{1}$. We discover that $\hat{D}_{1}$ can be treated as a biprojective bundle $\mathbb{P}(\mathcal{Q})\times_{\mathbb{P}^{4}}\mathbb{P}(\mathcal{Q})$. In addition, we interpret $\widetilde{D}_{1}$ as the blowup of $\hat{D}_{1}$ along the fibrewise diagonal $\hat{\Delta}_{D}$. We establish the structure $\mathsf{A}(\widetilde{D}_{1})$ by realizing that $\hat{\Delta}_{D}$ is isomorphic a flag variety and the normal bundle of $\hat{\Delta}_{D}$ in $\hat{D}_{1}$ is isomorphic to a subbundle of $\mathcal{T}_{\hat{\Delta}_{D}}$. With these tools, we manage to compute the excess intersection term. Then we assert that a curve $X\subseteq\mathbb{P}^{4}$ is algebraically skew if and only if the intersection class $[D_{1}^{\dagger}][\Gamma^{\dagger}]$ agrees with the excess intersection class $\varepsilon_{\Delta^{\dagger}_{\Gamma}}$ from the diagonal and conclude that the only algebraically skew curves in $\mathbb{P}^{4}$ are the rational normal curve and the elliptic normal curves.
@@ Acknowledgment
This project is partially supported by the NSF Award DMS-2142966. The author is a graduate student at Pennsylvania State University. The author is grateful to his advisors Sergei Tabachnikov and John Lesieutre for the weekly meetings in writing this paper. The author thanks Jack Huizenga for several crucial conversations and advice. The author is grateful to Edoardo Ballico for pointing out the relation between skew embeddings and Terracini loci and for Ciro Ciliberto for pointing out his work on the case of Terracini loci of curves in $\mathbb{P}^{4}$. The author also benefits from several insightful email exchanges with Joseph Landsberg, Serge Lvovski, and Igor Dolgachev. The author also thanks Brendan Hassett for pointing out the potential double point formula approach to the problem. The author also thanks Dawei Chen for a discussion of whether the intersection component on the blown-up diagonal in the present paper is a ribbon.
2. The basic bounds of $\mathsf{msdim}\,X$
In this section, we explain the existence of skew embeddings of a projective variety $X$ by observing the skewness of Veronese embeddings of degree $\geq 3$. We establish the general upper bound of the minimal complex algebraically skew embedding dimension $\mathsf{msdim}\,X$ by investigating the tangent of the secant variety. Then we introduce the degeneracy filtration of $\mathsf{Gr}(n,N)\times\mathsf{Gr}(n,N)$, and use them to compute a lower bound of $\mathsf{msdim}\,X$.
2.1. Skewness of higher degree Veronese embeddings
We first recall that rational curves of degree $\geq 3$ are algebraically skew. This follows from [BC21] or [BC23]. One can also prove this with straightforward computation.
@@ 2.1. Skewness of higher degree Veronese embeddings
By the observation above, one can prove that the following projections from degree $\geq 3$ Veronese embeddings are skew.
@@
(1) $\alpha\widetilde{\iota}_{d}(\widetilde{x})+\sum_{k=0}^{n}a_{k}\frac{\partial\widetilde{\iota}_{d}}{\partial x_{k}}(\widetilde{x})=\beta\widetilde{\iota}_{d}(\widetilde{y})+\sum_{k=0}^{n}b_{k}\frac{\partial\widetilde{\iota}_{d}}{\partial x_{k}}(\widetilde{y}).$
@@
2.2. The basic upper bound of $\mathsf{msdim}\,X$
To find the basic upper bound of $\mathsf{msdim}\,X$, suppose that $X\hookrightarrow\mathbb{P}^{N}$ is a skew embedding with $N=\mathsf{msdim}\,X$. Let $\mathsf{Sec}\,X\subseteq\mathbb{P}^{N}$ be the secant variety of $X$. By the Terracini’s Lemma [Ter11], see also [Zak05, Proposition II.1.10], we have